OpenStudy (anonymous):

the sum of hte three smallest of four consecutive oss integers is 30 greaterthan the largest integer. What is the largest integer?

OpenStudy (mertsj):

The integers are x x+1 x+2 x+3 x+(x+1)+(x+2)=(x+3)+30

OpenStudy (mertsj):

Can you solve that?

OpenStudy (anonymous):

is the beginning all added togather or is any of it multiplied?

OpenStudy (mertsj):

All added together.

myininaya (myininaya):

it says sum so that means we add (not multiply)

OpenStudy (mertsj):

did you get it?

OpenStudy (anonymous):

ya I got 10/3?

OpenStudy (mertsj):

But that is not an integer.

OpenStudy (anonymous):

?

myininaya (myininaya):

\[x+x+x+1+2=x+3+30\] \[3x+3=x+33\] try solving it from here

OpenStudy (mertsj):

Watch. I'll show you.

OpenStudy (mertsj):

too late. Myininaya will show you.

OpenStudy (anonymous):

x=15

myininaya (myininaya):

yep! :)

myininaya (myininaya):

now that is the smallest integer

myininaya (myininaya):

you are looking for the largest intger in your problem

OpenStudy (anonymous):

yes soo do I plug the 15 into the original equation?

myininaya (myininaya):

remember we had x <-smallest x+1 x+2 x+3 <- largest

myininaya (myininaya):

so if x=15 then x+3= what?

OpenStudy (anonymous):

18?

myininaya (myininaya):

thats right

OpenStudy (anonymous):

Great Thankyou!!!

OpenStudy (anonymous):

I understand a little better now